OpenStudy (mathmath333):

how many pairs of natural numbers are there the difference of whose squares is 45?

OpenStudy (ikram002p):

\(a^2-b^2=45\) hmm lets see

OpenStudy (gorv):

x^2-y^2=45

OpenStudy (gorv):

(x+y)(x-y)=45

OpenStudy (ikram002p):

ohhh lol its hyperbola

OpenStudy (gorv):

now possible factors of 45 are 1,3,5,9,15

OpenStudy (gorv):

9 and 6 7 and 2 or 22 and 23 are such numbers only

OpenStudy (gorv):

so only 3 are there

OpenStudy (mathmath333):

ok ,but is there algebraic way to solve it ? other than trial an error

OpenStudy (ikram002p):

its not trial its algebraic already

OpenStudy (gorv):

a^2 -b^2 = 45 (a+b)(a-b) = 45 45 = 45*1 or 9*5 or 15*3 if a+b =45 and a-b = 1 a = 23 and b = 22 if a+b = 9 and a-b = 5 a = 7 and b = 2 if a+b =15 and a-b = 3 a= 9 and b = 6

OpenStudy (mathmath333):

@gorv how did u find them

OpenStudy (gorv):

this the possible solution and we are considering each one by one

OpenStudy (gorv):

its not trial and error

OpenStudy (gorv):

(x+y)(x-y)=45 means 45 is multiplication of two numbers x+y and x-y that means its factor needs to be put in such way that they give product of 45